OpenStudy (anonymous):

I am having difficulties with this problem----- Give the most general solutions to the equation: 2sinxcosx - sin(2x)cos(2x) = 0 a. Simplify the first expression using the double-angle identity for sine. b. Factor the left side of the equation. c. Solve the factored equation.

OpenStudy (anonymous):

\[2\sin x \cos x-\sin2x \cos2x=2\sin x \cos x- 2\sin x \cos x(1-2\sin ^{2} x)=2\sin x \cos x-2\sin x \cos x+4\sin ^{3}x \cos x=4\sin ^{3} x \cos x\] so, \[4\sin ^{3} x \cos x=0\] so we get, \[\sin x=0, \cos x=0\] for \[\sin x=0,\] we get \[x=n \pi \] for \[n \in Z\] for \[\cos x=0, x= (n+1/2) \pi, n \in Z\]

OpenStudy (anonymous):

first equation is not visible fully, it will be: \[2 \sin x \cos x- 2 \sin x \cos x+ 4 \sin ^{3} x \cos x= 4 \sin ^ {3} x \cos x\]